OpenStudy (hellokitty17):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 2,500n - 4,000 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (hellokitty17):

@HootHoot

OpenStudy (hoothoot):

ok im trying this one throwing me off a bit

OpenStudy (hellokitty17):

ok! thanks

OpenStudy (hoothoot):

one year ago \[\Large\rm P(n)=-250n^2 + 2500n - 4000\]So the find `zeros`, we set the P(n) equal to zero.\[\Large\rm 0=-250n^2 + 2500n - 4000\]This will be easier to work with if we have no fancy coefficient on the squared term. So let's factor a -250 out of each term:\[\Large\rm 0=-250(n^2-10n+16)\]Then divide the -250 away,\[\Large\rm 0=n^2-10n+16\]It looks like this will factor nicely! Confused by any of those steps?

OpenStudy (hoothoot):

few dont wanna write that back down again

OpenStudy (hellokitty17):

Witch part is witch

OpenStudy (hellokitty17):

@HootHoot

oregonduck (oregonduck):

which one is part a b and c?

OpenStudy (hoothoot):

the mid is and the lower part is c i go from top to bottom

OpenStudy (hoothoot):

orry it is a lil confusing

OpenStudy (hoothoot):

but that was the best way to explain it the other ways made it harder to understand

OpenStudy (hellokitty17):

could you writ is again please with it saying part a b and c

oregonduck (oregonduck):

-250n^2 + 2500n - 4000 = 0 n^2 - 10n + 16 = 0 (n - 8)(n - 2) = 0 The zeroes are (0,2) and (0,8). The values between these two points are places where the profit is positive. At these two points, the promoter breaks even. *** n^2 - 10n + 16 = 0 -10/2 = -5 (-5)^2 = 25 n^2 - 10n + 25 = 9 (n - 5)^2 = 9 The maximum is when n - 5 = 0, or n = 5. P(5) = 9 * 250 = 2250, or (5,2250). *** The access of symmetry is at n = 5.

OpenStudy (hoothoot):

-250n^2 + 2500n - 4000 = 0 n^2 - 10n + 16 = 0 (n - 8)(n - 2) = 0 The zeroes are (0,2) and (0,8). The values between these two points are places where the profit is positive. At these two points, the promoter breaks even. n^2 - 10n + 16 = 0 -10/2 = -5 (-5)^2 = 25 n^2 - 10n + 25 = 9 (n - 5)^2 = 9 The maximum is when n - 5 = 0, or n = 5. P(5) = 9 * 250 = 2250, or (5,2250). The access of symmetry is at n = 5. ok i tried my best hope this works

OpenStudy (hoothoot):

my finger gett sore from typing